# Accurate stochastic simulation algorithm for multiscale models of infectious diseases

### v1

@@ Abstract
In the infectious disease literature, significant effort has been devoted to studying dynamics at a single scale. For example, compartmental models describing population-level dynamics are often formulated using differential equations. In cases where small numbers or noise play a crucial role, these differential equations are replaced with memoryless Markovian models, where discrete individuals can be members of a compartment and transition stochastically. Classic stochastic simulation algorithms, such as Gillespie’s algorithm and the next reaction method, can be employed to solve these Markovian models exactly. The intricate coupling between models at different scales underscores the importance of multiscale modelling in infectious diseases. However, several computational challenges arise when the multiscale model becomes non-Markovian. In this paper, we address these challenges by developing a novel exact stochastic simulation algorithm. We apply it to a showcase multiscale system where all individuals share the same deterministic within-host model while the population-level dynamics are governed by a stochastic formulation. We demonstrate that as long as the within-host information is harvested at a reasonable resolution, the novel algorithm we develop will always be accurate. Moreover, the novel algorithm we develop is general and can be easily applied to other multiscale models in (or outside) the realm of infectious diseases.
@@ 1 Introduction
Exact and approximate SSAs are derived from Gillespie’s work on models of molecular biology [Gil77]. For example, variants such as the Next Reaction Method [GB00] make clever use of the memorylessness of exponentially-distributed wait times to improve simulation times. On the other hand approximate methods such as Tau-Leaping [CGP06] differ from exact SSAs by implementing a bundle of events over a discrete time interval all at once, without updating the state of the system for each event, under the assumption that each individual event has only a small effect on propensities and that the changing of propensities over time can be approximated by state changes due to the cumulative effect of a bundle of events. With a few exceptions, SSAs are almost always applied under the Markovian assumption that $\alpha$ is (piecewise) constant between updates and $\tau$ is an exponentially-distributed wait time. However, in the case of infectious diseases, this is often not appropriate. For example, in situations where it is necessary to couple rapid events (such as events associated with within-host models) and relatively slow events (such as population-level changes), it becomes too computationally intensive to simulate all of the within-host events for each host in a population whilst simultaneously accounting for the evolution of behaviour at the population level in response to within-host temporal dynamics for each individual. Those events, in addition, are not Markovian in nature. Hence, there is a need for innovative methods which allow for accurate simulation in cases that involve multiple scales, and we address this problem in the current paper.
@@
In this paper, we present a novel simulation framework for multiscale infectious disease models. In Section 2, we develop the multiscale model and present the simulation algorithm. The simulation algorithm is the key innovation of this paper, as it exploits common features among individuals to deliver accurate and efficient results. The innovation of the algorithm allows for its adoption in a broad range of multiscale infectious disease models. Then in Section 3, we present simulation results for some key test problems which demonstrate the utility of the new algorithm. Finally, in Section 4, we discuss the implications of our new framework and simulation algorithm in a broader context.
@@ 2.1.1 Population-level model
The SI model (1) has tractable fixed points $(S,I)=(S^{*},I^{*})$. They are $(S^{*},I^{*})=(S_{0}^{*},I_{0}^{*})=(\Lambda\mu^{-1},0)$ and $(S^{*},I^{*})=(S_{1}^{*},I_{1}^{*})=(\mu\beta^{-1},[\Lambda\beta-\mu^{2}](\beta\mu)^{-1})$. We shall refer to $(S_{0}^{*},I_{0}^{*})$ as the elimination state and $(S_{1}^{*},I_{1}^{*})$ as the endemic state. It is possible to show through linear stability and bifurcation analysis that a transcritical bifurcation occurs when $\Lambda\beta=\mu^{2}$. If $\Lambda\beta>\mu^{2}$, then the disease becomes endemic with $(S_{1}^{*},I_{1}^{*})$ being positive and stable whilst the disease elimination state $(S_{0}^{*},I_{0}^{*})$ is unstable. If $\Lambda\beta<\mu^{2}$, then $(S_{1}^{*},I_{1}^{*})$ has a negative infectious population and is therefore an irrelevant fixed point and $(S_{0}^{*},I_{0}^{*})$ is stable.
@@
    Use a stochastic simulation algorithm (SSA) at the population level. We note that as an ODE, the model only describes large continuous populations, and when populations $S$ or $I$ are small, the model is inappropriate. For example, if $I_{0}>0$ and $\Lambda\beta>\mu^{2}$, then $I$ may get arbitrary small. However, since $(S_{0}^{*},I_{0}^{*})$ is unstable, disease elimination is impossible. On the other hand, even if $\Lambda\beta>\mu^{2}$, elimination should be possible when $I$ is small. Because in reality a small number of infected individuals may, stochastically, recover or die before infecting anybody. Such an elimination event can be achieved by a stochastic simulation of model (1). For instance, a discrete and stochastic transition of $I=1$ individual to $I=0$ leads exactly to the elimination steady state $(S_{0}^{*},I_{0}^{*})$. This is an absorbing state since small perturbations are not possible in the discrete stochastic (and thus realistic) scenario. We shall demonstrate in Subsection 2.1.2 how to extend model (1) into a stochastic simulation algorithm. Although for model (1), such conversion is widely understood, we shall highlight particular ideas useful for generating a multiscale model.
@@
    Introduce a deterministic within-host model for individual infections. We note that one big assumption in model (1) is that the rates (or propensities) for population-level events to take place do not evolve during the inter-event time intervals. However, such an assumption may be an oversimplification in a real-world scenario. For example, each individual’s mortality rate at the population level can be related to its within-host pathogen load and immune responses [GMN14]. Moreover, a variety of experimental observations [EGB12, BWE+14] indicate that temporal changes in an infectious agent’s within-host viral load can be associated with its propensity to infect others. We, therefore, extend our model (1) by introducing a deterministic within-host model for the infectious population. In Subsection 2.1.2, we will highlight how time-dependent event propensities complicate the SSA algorithm. Then in Subsection 2.1.3, we will present a simple deterministic within-host model.
@@ 2.1.2 Converting population-level model into an SSA
The population-level model shown in (1) describes a system of deterministic interactions that govern the dynamics of an infectious disease. A key assumption behind the ODE formualtion is that populations are large and continuous, and hence noise can be ignored. The underlying discrete stochastic process of (1) is a continuous time Markov chain (CTMC), which describes the evolution of the discrete stochastic non-negative state $(\mathcal{S}(t),\mathcal{I}(t))\in\mathbb{N}_{\geq 0}^{2}$. $\mathcal{S}(t)$ and $\mathcal{I}(t)$, respectively, represent the number of susceptible and infectious individuals at time $t\in[0,\infty)$. They are piece-wise constant functions in time because the system remains in a constant state for an exponentially distributed amount of time until moving to the next state [All10]. Moreover, a CTMC has the Markovian property of being memory-less; its future state only depends on the current state of the system [All17].
Consider an arbitrary time $t$ when $\mathcal{S}(t)=s$ and $\mathcal{I}(t)=i$, we the define transition probabilities of the stochastic process from state $(s,i)$ to state $(s+m,i+n)$ by time $t+\Delta t$ as
@@
The stochastic simulations of (3) – exact or approximate – are well-known in the literature and can be achieved in a number of ways. The simplest approach uses a small set time step $\Delta t$, thus it is referred to as a time-driven algorithm. At each moment in time of a simulation, the current state of the system $(\mathcal{S}(t),\mathcal{I}(t))=(s,i)$ is known. Then, multiplying (3) by $\Delta t$ gives the update formula to the leading order:
@@
Two exact methods, the Gillespie algorithm (direct SSA) and Next Reaction Method (NRM) solve this problem by asking the question
@@
The use of the upper case $T$ emphasises that (10) is a distribution for a wait time from the current time $t$ (i.e. for a putative calendar time of $t+T$). The putative (event wait) time, $\tau$ can be drawn from $\tau\sim\phi_{(s,i)}(T)$. The use of the term ‘putative’ is common. This is because when $\tau$ is sampled, the event is scheduled for the calendar time $t+\tau$. But this schedule may change if something changes the propenensity $\alpha(s,i)$. In the direct SSA, all events are considered under a single propensity, so the putative time will always result in an event at $t+\tau$. However, this is not the case for methods like the NRM.
The direct SSA and NRM take theoretically identical but practically different approaches to the question in (9). In the direct SSA, inverse transform sampling is used to draw an event time using the distribution (10):
$\tau=\alpha(s,i)^{-1}\log\left(u^{-1}\right)=\Phi^{-1}(1-u),$ (11)
where $u\sim U(0,1)$ and $\Phi$ is the cumulative distribution function associated with the density $\phi$. After determining when an event occurs, the direct SSA then determines which event occurs based on the relative sizes of the probabilities (5)–(8). For example, the probability of a birth event (i.e. $(s,i)\rightarrow(s+1,i)$) at calendar time $t+\tau$ is $\alpha^{(b)}/\alpha(s,i)$, similarly for the other events in proportion to their respective propensities. However for NRM, instead of determining when an event occurs followed by which event occurs, it reverses the order of these calculations by making use of two special properties of the exponential distribution (10). Firstly, given any sensible partition of the state propensity by event type $\alpha=\sum_{(e)\in\mathcal{E}}\alpha^{(e)}$ where $\mathcal{E}$ are the set of possible events (e.g., in this manuscript we have that =$\mathcal{E}=\{(b),(ds),(di),(i)\}$), (11) has the same probability density function as
@@
where $u^{(e)}\sim U(0,1)$ is independently sampled for each possible event and $\tau^{(e)}$ is the putative time associated with realising each event type. The most imminent event $(e)$ – the one with the smallest putative wait time – is the event which is chosen to propagate the simulation forward. Importantly, because of the exponential distribution (10), the event corresponding to this most imminent time has a probability that matches the probability of the event determined in the direct SSA. Therefore in the NRM, putative times for event types ($\tau^{(e)}=(\alpha^{(e)})^{-1}\log((z^{(e)})^{-1})$) are drawn and stored as an array. The minimum event putative time ($\tau=\tau^{(m)}$, $(m)\in\mathcal{E}$) is then determined to be the ‘next reaction’ and the state is updated according to the event $(m)$ associated with this putative time. Interestingly, the second property of the exponential distribution (10) then allows for the NRM to be efficient. Instead of throwing away and resampling (rescheduling) all putative times for all events $(e)\neq(m)$, the times $\tau^{(e)}$ are kept and updated by first subtracting $\tau^{(m)}$ ($\tau^{(e)^{\prime}}=\tau^{(e)}-\tau^{(m)})$. We use the dash notation to indicate that a wait time has been updated to a new wait time due to the moving forward of time. Then the state is changed based on the event $(m)$, which subsequently changes many of the event propensities. Because each of the putative times $\tau^{(e)}$ have been initially sampled from exponential distributions, their updated values $\tau^{(e)^{\prime}}$ are conditioned on $\tau^{(e)}>\tau^{(m)}$, which are also exponentially distributed random numbers with propensity $\alpha^{(e)}$ (the lack of markings here highlighting that this is the propensity prior to any state change due to event $(m)$) and there is no need to resample the putative event times which are computationally expensive to sample. An updated putative time needs to be generated for the event $(m)$. However it is clear from the sampling equation (11) that all other events $(e)\neq(m)$ just need to be scaled after the event $(m)$:
@@
where the bar indicates that the quantity has been updated for the change of current time from $t$ to $t+\tau^{(m)}$ (the dashed notation), and also updated for the change of state/propensity associated with event $(m)$ from the time $t+\tau^{(m)}$ to the new putative calendar event time of $t+\tau^{(m)}+\bar{\tau}$. After all calculations are done to update the state and putative times for the event $(m)$, $t+\tau^{(m)}$ becomes the current time $t$ and the new state becomes the current state (so bars and dashes are subsequently dropped to be used to denote the updates for the next event). Using the Gillespie direct SSA or the NRM, an exact stochastic simulation of the model is possible. In 1(b) we show simulations of this algorithm against the deterministic ODE model (1).
@@
Figure 1: (a) Schematic diagram for the population-level model. The boxes represent different compartments while the arrows denote transition between compartments. The compartment linked to the tail end of the arrow, ()maps-to($\mapsto$)( ↦ ), does not experience any population flow in the corresponding transition. The dashed arrows denote the removal rates. (b) $\mathbf{100}$ realisations of the population-level model. The stochastic simulations are implemented based on Gillespie’s algorithm while the deterministic result is generated by MATLAB’s $ode15s$ solver with both absolute and relative tolerances being $10^{-8}$. We set the initial condition $\left(S_{0},I_{0}\right)$ as $(80,20)$ people, and the ending time of the simulation to be $50$ days. The parameter values chosen are: $\Lambda=10^{-3}\text{ person}\times\text{day}^{-1}$, $\beta=1.5\times 10^{-3}\rm\text{ }day^{-1}\rm person^{-1}$, and $\mu=10^{-5}\rm\text{ }day^{-1}$. A similar result can be obtained using the NRM.
@@
The purpose of discussing these known stochastic methods in detail here is to highlight the importance of the Markovian property in the population-level system. That is, event times are exponentially distributed (10) and do not depend on history or time explicitly between events (when the state of the system is constant). One nontrivial issue, which we shall address in detail, can arise if the state of the system is changing in time between events. In multiscale (for example within-host and population-level coupled) models, population-level propensities can depend on within-host dynamics, and thus change with time between population-level state changes. The obvious solution here is to attempt to adapt the within-host model into the SSA and generate putative times for all events (including those within each infectious individual). In this respect, all events, irrespective of whether they are population-level or within-host level would be modelled exactly. However, it is completely infeasible to do this since: 1) there may be a large number of infectious individuals; 2) each individual will have their own SSA; and 3) there may be an extremely large number of within-host events simulated before a single population-level model event (the multiscale problem). Instead, because the within-host dynamics are associated with much smaller timescales compared to the population-level dynamics, it becomes crucial to develop an appropriate method for running an SSA at the population level subject to deterministic and smooth time-varying propensities (in our case as a result of changing viral load of each individual). We are therefore faced with the following two challenges: 1) time varying propensities: propensities that may change with time between discrete event times; and 2) the individuality problem: temporal change in propensities may be complex due to individual contributions and subject to historical events (for example the previous times when individuals become infected). The time varying propensity problem has known solutions and can be solved exactly as we shall demonstrate. Of course, sometimes the additional work required to deal with this problem exactly is not considered to be worth the effort and approximate algorithms, such as $\tau$-leaping, are used (appropriate when the relative change in propensity between events is sufficiently small and therefore approximated to be constant) [Gil01]. The individuality problem is a much harder problem. This is unsurprising since more detail in SSA simulations and, indeed, any model, will always increase computational demands. To overcome these two challenges, we introduce in the next section a simple within-host model to use as a test problem and we assume that, once infected, all individuals deterministically trace solutions of this model. This means that what separates individuals is the initial time of infection only.

### v2

@@ Abstract
In the infectious disease literature, significant effort has been devoted to studying dynamics at a single scale. For example, compartmental models describing population-level dynamics are often formulated using differential equations. In cases where small numbers or noise play a crucial role, these differential equations are replaced with memoryless Markovian models, where discrete individuals can be members of a compartment and transition stochastically. Classic stochastic simulation algorithms, such as the next reaction method, can be employed to solve these Markovian models exactly. The intricate coupling between models at different scales underscores the importance of multiscale modelling in infectious diseases. However, several computational challenges arise when the multiscale model becomes non-Markovian. In this paper, we address these challenges by developing a novel exact stochastic simulation algorithm. We apply it to a showcase multiscale system where all individuals share the same deterministic within-host model while the population-level dynamics are governed by a stochastic formulation. We demonstrate that as long as the within-host information is harvested at a reasonable resolution, the novel algorithm will always be accurate. Furthermore, our implementation is still efficient even at finer resolutions. Beyond infectious disease modelling, the algorithm is widely applicable to other multiscale systems, providing a versatile, accurate, and computationally efficient framework.
@@ 1 Introduction
Exact and approximate SSAs are derived from Gillespie’s foundational work on models of molecular biology [Gil77]. For instance, variants such as the Next Reaction Method [GB00] make clever use of the memorylessness of exponentially-distributed wait times to improve simulation times. On the other hand, approximate methods such as Tau-Leaping [CGP06] allow for significant computational savings by implementing a bundle of events over a discrete time interval all at once. It forgoes updating the system state after every individual event, under the assumptions that (1) each individual event has only a small effect on propensities and (2) the changing of propensities over time can be approximated by state changes due to the cumulative effect of a bundle of events. With a few exceptions, SSAs are almost always applied under the Markovian assumption that the event propensities $\alpha$ is (piecewise) constant between updates and $\tau$ is an exponentially-distributed wait time. However, in the case of infectious diseases, this is often not appropriate. For example, in situations where it is necessary to couple rapid events (such as events associated with within-host models) and relatively slow events (such as population-level changes), it becomes too computationally intensive to simulate all of the within-host events for each host in a population whilst simultaneously accounting for the evolution of behaviour at the population level in response to within-host temporal dynamics for each individual. Those events, in addition, are not Markovian in nature, so that event propensities $\alpha$ become functions of system states and time and the distribution of $\tau$ is much more complex. Hence, there is a need for innovative methods which allow for accurate simulation in cases that involve multiple scales. We investigate this problem by developing an exact simulation framework for multiscale infectious disease models that incorporate non-Markovian dynamics.
Previous efforts have been made under the assumption that the distribution of event wait times, $\tau$, is known, analytic, and fixed. For example, an approximate algorithm based on series expansion [BLTS14], exact algorithms using the Laplace transform [MR18] and based on next reaction method [GB00], and an algorithm for non-Markovian processes on time-varying networks [VG15] have been proposed. In contrast, our work relaxes this assumption, allowing $\tau$ to follow an arbitrary time-varying distribution as defined by the multiscale infectious disease model. Moreover, our work incorporates individuality by considering how each individual’s within-host system influences $\tau$. Researchers have also employed a second-order series expansion approximation to avoid the computational expense of numerically integrating the propensity function $\alpha$ in non-Markovian settings [CF12]. This method assumes (1) $\alpha$ and its derivatives have well-defined analytic forms at the population level, and (2) the time variation in $\alpha$ is sufficiently slow for a second-order approximation to be accurate. However, these assumptions may not hold in multiscale infectious disease systems. For example, if disease transmission depends on each individual’s within-host viral load, then the transmission propensity $\alpha$ will vary for each individual, and the functional form of $\alpha$ may not be analytically expressed. Additionally, if the within-host dynamics evolve rapidly, a second-order approximation may not suffice. Our work addresses these limitations by relaxing these constraints, thereby enhancing the generality of our framework.
Another relevant algorithm, the ‘modified next reaction method’, was introduced by Anderson [And07]. In this approach, the wait time $\tau$ for an event is calculated by taking the antiderivative of the propensity function $\alpha$, which depends on both the system state and time, and then performing inversions. The primary distinction between this method and our novel framework lies in the definition of $\alpha$. Instead of defining $\alpha$ as a function that varies with the system state and time, we define $\alpha$ to be state-invariant by leveraging information about how infectious disease models at different scales couple to one another at the level of each infected individual. This formulation enables us to compute the integrated propensity function only once for the entire simulation. We then store the computed values in look-up tables for efficient inversion, incorporating resampling and rescaling of the wait time $\tau$. Hence, another key contribution of our work is its enhanced practicality.
@@
The paper is structured as follows. In Section 2, we develop the multiscale model and present the simulation algorithm. The simulation algorithm is the key innovation of this paper, as it exploits common features among individuals to deliver accurate and efficient results. The innovation of the algorithm allows for its adoption in a broad range of multiscale infectious disease models. Then in Section 3, we present simulation results for some key test problems that highlight the utility of the new algorithm. We show that our algorithm achieves superior accuracy without sacrificing efficiency. Finally, in Section 4, we discuss the implications of our new framework and simulation algorithm in a broader context.
@@ 2.1.1 Population-level model
    Use a stochastic simulation algorithm (SSA) at the population level. We note that as an ODE, the model only describes large continuous populations, and when populations $S$ or $I$ are small, the model is inappropriate (see Section 7.1 in Supplementary Information). We will demonstrate in Section 2.1.2 how to extend model (1) into a stochastic simulation algorithm. Although for model (1), such conversion is widely understood, we shall highlight particular ideas useful for generating a multiscale model.
@@
    Introduce a deterministic within-host model for individual infections. We note that one big assumption in model (1) is that the rates (or propensities) for population-level events to take place do not evolve during the inter-event time intervals. However, such an assumption may be an oversimplification in a real-world scenario. A variety of experimental observations [EGB12, BWE+14] indicate that temporal changes in an infectious agent’s within-host viral load can be associated with its propensity to infect others. We, therefore, extend our model (1) by introducing a deterministic within-host model for the infectious population. In Section 2.1.2, we will highlight how time-dependent event propensities complicate the SSA algorithm. Then in Section 2.1.3, we will present a simple deterministic within-host model.
@@ 2.1.2 Converting population-level model into an SSA
The underlying discrete stochastic process of population-level model (1) is a continuous time Markov chain (CTMC), which describes the evolution of the discrete stochastic non-negative state $(\mathcal{S}(t),\mathcal{I}(t))\in\mathbb{N}_{\geq 0}^{2}$. $\mathcal{S}(t)$ and $\mathcal{I}(t)$, respectively, represent the number of susceptible and infectious individuals at time $t\in[0,\infty)$. They are piece-wise constant functions in time because the system remains in a constant state for an exponentially distributed amount of time until moving to the next state [All10]. Moreover, a CTMC has the Markovian property of being memory-less; its future state only depends on the current state of the system [All17].
Consider an arbitrary time $t$ when $\mathcal{S}(t)=s$ and $\mathcal{I}(t)=i$, we then define transition probabilities of the stochastic process from state $(s,i)$ to state $(s+m,i+n)$ by time $t+\Delta t$ as
@@
The stochastic simulations of (3) – exact or approximate – are well-known in the literature and can be achieved in a number of ways. The simplest approximate approach uses a small set time step $\Delta t$, thus it is referred to as a time-driven algorithm. At each moment in time of a simulation, the current state of the system $(\mathcal{S}(t),\mathcal{I}(t))=(s,i)$ is known. Then, multiplying (3) by $\Delta t$ gives the update formula to the leading order:
@@
We then introduce the idea underlying the exact Next Reaction Method (NRM), as we will extend and apply a similar strategy to a later multiscale system in Section 2.2.2. We consider the forward Kolmogorov equation and ask the question
@@
The use of the upper case $T$ emphasises that (10) is a distribution for a wait time from the current time $t$ (i.e. for a putative calendar time of $t+T$). The putative (event wait) time, $\tau$ can be drawn from $\tau\sim\phi_{(s,i)}(T)$. The use of the term ‘putative’ is common. This is because when $\tau$ is sampled, the event is scheduled for the calendar time $t+\tau$. Note this schedule may change if something changes the propensity $\alpha(s,i)$.
The strategy employed in the NRM to address the question in (9) relies on two key properties of the exponential distribution (10). First, given any sensible partition of the state propensity by event type, $\alpha=\sum_{(e)\in\mathcal{E}}\alpha^{(e)}$, where $\mathcal{E}$ represents the set of possible events (e.g., in this manuscript, =$\mathcal{E}=\{(b),(ds),(di),(i)\}$), inverse sampling can be utilised to determine the next event time:
@@
where $u^{(e)}\sim U(0,1)$ is independently sampled for each possible event and $\tau^{(e)}$ is the putative time associated with realising each event type. The most imminent event, denoted as event $(m)$, is the one with the smallest putative wait time and is chosen to propagate the simulation forward (i.e., the “next reaction”).
The second property of the exponential distribution (10) enables the NRM to update putative times when propensities change without discarding and resampling (rescheduling) putative times for all events $(e)\neq(m)$, recall $(m)$ is the most imminent event. Instead, the times $\tau^{(e)}$ are retained and updated by first subtracting $\tau^{(m)}$ ($\tau^{(e)’}=\tau^{(e)}-\tau^{(m)}$). We use the dash notation to indicate that a wait time has been updated to a new wait time due to the moving forward of time. Then the state is changed based on the event $(m)$, which subsequently changes many of the event propensities. Because each of the putative times $\tau^{(e)}$ have been initially sampled from exponential distributions, their updated values $\tau^{(e)^{\prime}}$ are conditioned on $\tau^{(e)}>\tau^{(m)}$, which are also exponentially distributed random numbers with propensity $\alpha^{(e)}$ (the lack of markings here highlighting that this is the propensity prior to any state change due to event $(m)$) and there is no need to resample the putative event time. Thus, a new putative time needs to be generated for the event $(m)$. Moreover, the putative times for all other events $(e)\neq(m)$ only need to be rescaled following the occurrence of event $(m)$, as shown in (12):
@@
where the bar indicates that the quantity has been updated for the change of current time from $t$ to $t+\tau^{(m)}$ (the dashed notation), and also updated for the change of state/propensity associated with event $(m)$ from the time $t+\tau^{(m)}$ to the new putative calendar event time of $t+\tau^{(m)}+\bar{\tau}$. After all calculations are done to update the state and putative times for the event $(m)$, $t+\tau^{(m)}$ becomes the current time $t$ and the new state becomes the current state (so bars and dashes are subsequently dropped to be used to denote the updates for the next event). In 1(b) we show simulations of this algorithm against the deterministic ODE model (1).
@@
Figure 1: (a) Schematic diagram for the population-level model. The boxes represent different compartments while the arrows denote transition between compartments. The compartment linked to the tail end of the arrow, ()maps-to($\mapsto$)( ↦ ), does not experience any population flow in the corresponding transition. The dashed arrows denote the removal rates. (b) $\mathbf{100}$ realisations of the population-level model. The stochastic simulations are implemented based on the strategy of NRM while the deterministic result is generated by MATLAB’s $ode15s$ solver with both absolute and relative tolerances being $10^{-8}$. We set the initial condition $\left(S_{0},I_{0}\right)$ as $(80,20)$ people, and the ending time of the simulation to be $50$ days. The parameter values chosen are: $\Lambda=10^{-3}\text{ person}\times\text{day}^{-1}$, $\beta=1.5\times 10^{-3}\rm\text{ }day^{-1}\rm person^{-1}$, and $\mu=10^{-5}\rm\text{ }day^{-1}$.
@@
The purpose of discussing the strategy used in NRM in detail here is to highlight the importance of the Markovian property in the population-level system. That is, event times are exponentially distributed (10) and do not depend on history or time explicitly between events (when the state of the system is constant). One nontrivial issue, which we shall address in detail, can arise if the state of the system is changing in time between events. In multiscale (for example within-host and population-level coupled) models, population-level propensities can depend on within-host dynamics, and thus change with time between population-level state changes. The obvious solution here is to attempt to adapt the within-host model into the SSA and generate putative times for all events (including those within each infectious individual). In this respect, all events, irrespective of whether they are population-level or within-host level would be modelled exactly. However, it is completely infeasible to do this since: 1) there may be a large number of infectious individuals; 2) each individual will have their own SSA; and 3) there may be an extremely large number of within-host events simulated before a single population-level model event (the multiscale problem). Instead, because the within-host dynamics are associated with much smaller timescales compared to the population-level dynamics, it becomes crucial to develop an appropriate method for running an SSA at the population level subject to deterministic and smooth time-varying propensities (in our case as a result of changing viral load of each individual).
We are therefore faced with the following two challenges: 1) time varying propensities: propensities that may change with time between discrete event times; and 2) the individuality problem: temporal change in propensities may be complex due to individual contributions and subject to historical events (for example the previous times when individuals become infected). The time varying propensity problem has known solutions and can be solved exactly as we shall demonstrate. Of course, sometimes the additional work required to deal with this problem exactly is not considered to be worth the effort and approximate algorithms, such as $\tau$-leaping, are used (appropriate when the relative change in propensity between events is sufficiently small and therefore approximated to be constant) [Gil01]. The individuality problem is a much harder problem. This is unsurprising since more detail in SSA simulations and, indeed, any model, will always increase computational demands. To overcome these two challenges, we introduce in the next section a simple within-host model to use as a test problem and we assume that, once infected, all individuals deterministically trace solutions of this model. This means that what separates individuals is the initial time of infection only.